OpenStudy (anonymous):

Solve t = p+3r/6 for r.

jhonyy9 (jhonyy9):

how is correct ? t = (p+3r)/6 or ???

OpenStudy (jackellyn):

Multiply both sides by 6 6t=p+3r subtract p from both sides 6t-p=3r what would be the next step to get r alone?

OpenStudy (anonymous):

P+3r is over 6

OpenStudy (anonymous):

t=p+3r/6 \[\frac{p+3r}{6}(6)=t\] p+3r=6t -p -p ________________ 3r=6t-p divide 3 on bot sides r=6t-p/3

OpenStudy (anonymous):

r = p/3 + 3t